OpenStudy (anonymous):

integral(1/(1+cosx)) step by step please

myininaya (myininaya):

for some reason I'm thinking of multiplying by bottom's conjugate on top and bottom first

myininaya (myininaya):

This should give you a pretty easy form to integrate afterwards of course you might want to write the resulting fraction as two fractions

OpenStudy (anonymous):

that was my instinct also but it needs some fancy u sub after that in order work i think

myininaya (myininaya):

fancy sub not needed

myininaya (myininaya):

reg sub needed :p

OpenStudy (anonymous):

itll look like integral((cscx)^2 -csc(x)cot(x))

OpenStudy (anonymous):

then what

myininaya (myininaya):

well the derivative of cot(x) is -csc^2(x) and the derivative of csc(x) is -csc(x)cot(x) isn't ?

myininaya (myininaya):

isn't it?*

OpenStudy (anonymous):

im lost. so what would be the next step?

myininaya (myininaya):

if you know the derivative of cot(x) is -csc^2(x) then you know the integral of -csc^2(x) is cot(x)

myininaya (myininaya):

if you know the integral of -csc^2(x) is cot(x) then you know integral of csc^2(x) is -cot(x)

myininaya (myininaya):

well of course +c

OpenStudy (anonymous):

oh ya good point

OpenStudy (anonymous):

problem is wolfram says: tan(x/2)

OpenStudy (anonymous):

how do i convert it to that?

myininaya (myininaya):

so you see we have \[-\cot(x)+\csc(x)+C\] now: let's see tan(x/2)... \[\tan(x/2)=\frac{\sin(x)}{1+\cos(x)}\] What would happen if you wrote your answer in terms of sin and cos and combine the fractions

OpenStudy (anonymous):

(1-cosx)/sinx - no?

OpenStudy (anonymous):

oh got it

OpenStudy (anonymous):

that can be multiplied w. conjugate to sinx/(1+cosx) thanks

OpenStudy (anonymous):

thanks

myininaya (myininaya):

great stuff

OpenStudy (anonymous):

any other ways of doing this problem that might be quicker?

myininaya (myininaya):

well you weren't asked to show it was tan(x/2)+C were you just wanted to see if you got the same as wolfram right?

OpenStudy (anonymous):

ya

myininaya (myininaya):

Off the top of my head I can't think of a shorter way but actually I remember for some trig expressions or when integrating some trig expressions you might want to consider trig substitutions like u=tan(x/2) I think there might be another one let me see if I can find it

myininaya (myininaya):

http://www.math.ucsd.edu/~ebender/20B/trig.pdf I wouldn't exactly call it shorter though

OpenStudy (anonymous):

ok cool thanks